OpenStudy (anonymous):

Tania bought a glass pyramid from a gift store. The pyramid, with a height of 4 cm and a square base of side 6 cm, was packed in a box shaped as a cube of side 8 cm. The empty space in the box was packed with paper tissue. How many cubic centimeters of the box were stuffed with paper tissue? Answer 192 cubic centimeters 368 cubic centimeters 464 cubic centimeters 480 cubic centimeters

OpenStudy (hoblos):

volume of pyramid = (area of base)(height)/3 = (6*6)(4)/3 = 48 cm^3 volume of box = side^3 = 8^3 = 512 cm^3 volume of the part stuffed with paper tissue = volume of box - volume of pyramid so it is 512 - 48 = 464 cm^3

OpenStudy (anonymous):

thank-you so much ive got so much to get done today your help is greatly appreciated! =)

OpenStudy (hoblos):

glad to help dear =)